solve and check
5y-10+y=7y+18-5y

Answer by stanbon(75887)

solve and check
5y-10+y=7y+18-5y
----
6y - 10 = 2y + 18
4y = 28
y = 7
-----
Check:
5y-10+y=7y+18-5y
5*7 - 10 + 7 = 7*7 + 18 - 5*7
35 -3 = 49-17
32 = 32
=============
Cheers,
Stan H.